# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ A.2 Section 2
Take $\mathbb{A}=\{1,2\}$ and $\mathbb{T}=\{0,1\}$. Recall that $\Omega=\{\omega_{1},\omega_{2}\}$, with $\omega_{k}=k$ for $k\in\{1,2\}$, and $\mathscr{A}=\mathcal{P}(\Omega)$.
For the first example, we may set $W=\Omega\times\mathbb{A}^{\mathbb{T}}$. Any $(\omega,f)\in W$ can be identified with the triple $(\omega,f(0),f(1))$, so that $W$ corresponds to the set denoted by $W$ in the example in subsection 2.3. Under this identification, it is easily verified that $D_{t,f}=\Omega$ for all $f\in\mathbb{A}^{\mathbb{T}}$ and $t\in\mathbb{T}$, ${\bm{\mathsf{x}}}_{0}={\bm{\mathsf{x}}}_{0}(f)$ for all $f\in\mathbb{A}^{\mathbb{T}}$, and ${\bm{\mathsf{x}}}_{k}={\bm{\mathsf{x}}}_{1}(f)$ for all $f\in\mathbb{A}^{\mathbb{T}}$ with $f(0)=k$, for both $k=1,2$. Assumptions AP.W0 to AP.W3 are readily verified. Under this identification, the action path construction yields exactly the objects $F$, $\pi$, and ${\bm{\mathsf{X}}}$ from the example in subsection 2.3.
For the second example, let $W^{\prime}$ as in subsection 2.3. Interpret every triple $(\omega,k,m)\in W^{\prime}$ as the pair $(\omega,f)$ where $f\in\mathbb{A}^{\mathbb{T}}$ is given by $f(0)=k$ and $f(1)=m$. Interpret the pair $(\omega_{1},2)$ as the map $f\colon\mathbb{T}\to\{0,1,2\}$ with $f(0)=2$ and $f(1)=0$. Then, $W\subseteq{\mathbb{A}^{\prime}}^{\mathbb{T}}$ with $\mathbb{A}^{\prime}=\mathbb{A}\cup\{0\}$, where 00 is a placeholder for inaction. Furthermore, $D_{t,f}=\Omega$ for all $(t,f)\in\mathbb{T}\times{\mathbb{A}^{\prime}}^{\mathbb{T}}$ with $t=0$, or $t=1$ and $f(0)=1$; and $D_{1,f}=\{\omega_{2}\}$ for $f\colon\mathbb{T}\to\mathbb{A}^{\prime}$ with $f(0)=2$; for all other pairs $(t,f)\in\mathbb{T}\times{\mathbb{A}^{\prime}}^{\mathbb{T}}$ we have $D_{t,f}=\emptyset$. Moreover, ${\bm{\mathsf{x}}}^{\prime}_{0}={\bm{\mathsf{x}}}_{0}(f)$ for all $f\in{\mathbb{A}^{\prime}}^{\mathbb{T}}$; and ${\bm{\mathsf{x}}}^{\prime}_{1}={\bm{\mathsf{x}}}_{1}(f)$ for all $f\colon\mathbb{T}\to\mathbb{A}^{\prime}$ with $f(0)=1$; and ${\bm{\mathsf{x}}}^{\prime}_{2}={\bm{\mathsf{x}}}_{1}(f)$ for $f\colon\mathbb{T}\to\mathbb{A}^{\prime}$ with $f(0)=2$. Assumptions AP.W0 to AP.W3 are readily verified. Under these identifications, the action path construction yields exactly the objects $F^{\prime}$, $\pi^{\prime}$, and ${\bm{\mathsf{X}}}^{\prime}$ from the example in subsection 2.3. ∎
@@
Clearly, for all $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$, $D_{t,f}=\Omega$, whence AP.W0.
Regarding AP.W1, let $w=(\omega,f)\in W$ and $t,u\in\mathbb{T}$ with $t\neq u$ such that $x_{t}(w)=x_{u}(w)$. Without loss of generality, assume $t<u$. If $\mathbb{A}$ is a singleton, then $\mathbb{A}^{\mathbb{T}}$ is a singleton as well. Hence, $x_{t}(w)=\{w\}$. But $\mathbb{A}$ must be a singleton, since else there would be $g\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,g)\in x_{t}(w)$ and $g(t)\neq f(t)$. Thus, $(\omega,g)\in x_{t}(w)\setminus x_{u}(w)$ – a contradiction.
Regarding AP.W2, let $\omega\in\Omega$, $\tilde{f}\in\mathbb{A}^{\mathbb{T}}$, and $\mathbb{T}^{\prime}\subseteq\mathbb{T}$ satisfying $x_{t}(\omega,\tilde{f})\in F$ for all $t\in\mathbb{T}^{\prime}$. Then we already have $(\omega,\tilde{f})\in W$.
AP.W3 is clearly satisfied because $D_{t,f}=\Omega$ for all $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$.
(The timing problem):  Clearly, for any $\omega\in\Omega$ there is $f\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,f)\in W$, e.g., the constant map taking only the value $1$ at all times, in all components.
Regarding assumption AP.W0, for all $(t,f)\in\mathbb{A}^{\mathbb{T}}$, $D_{t,f}=\emptyset$ if $f$ is not decreasing on $[0,t)_{\mathbb{T}}$ or if $f(t-)=0$. Else, $D_{t,f}=\Omega$. Hence, AP.W0 is satisfied.
Regarding assumption AP.W1, let $w=(\omega,f)\in W$ and $t,u\in\mathbb{T}$ with $t<u$ such that $x_{t}(w)=x_{u}(w)$. Hence, for all decreasing $g\colon\mathbb{T}\to\mathbb{A}$ with $g|_{[0,t)_{\mathbb{T}}}=f|_{[0,t)_{\mathbb{T}}}$ we must have $f(v)=g(v)$ for all $v\in[t,u)_{\mathbb{T}}$. Thus, we must have $f(t-)=0$. Hence, $x_{t}(w)=\{(\omega,f)\}$.
Regarding assumption AP.W2, let $\omega\in\Omega$, $\tilde{f}\in\mathbb{A}^{\mathbb{T}}$, and $\mathbb{T}^{\prime}\subseteq\mathbb{T}$ such that $x_{t}(\omega,\tilde{f})\in F$ for all $t\in\mathbb{T}^{\prime}$. Hence, $\tilde{f}$ is decreasing on $[0,\sup\mathbb{T}^{\prime})_{\bar{\mathbb{R}}_{+}}\cap\mathbb{T}$ (where, in the context of the order on $\mathbb{T}$, we have $\sup\emptyset=0$). Hence, there is decreasing $f\colon\mathbb{T}\to\mathbb{A}$ such that $f|_{[0,t)_{\mathbb{T}}}=\tilde{f}|_{[0,t)_{\mathbb{T}}}$ for all $t\in\mathbb{T}^{\prime}$. By construction, $(\omega,f)\in W$.
Regarding assumption AP.W3, it is sufficient to note that $D_{t,f}$ equals $\emptyset$ or $\Omega$, for all $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$ which has been established earlier.
@@
Regarding assumption AP.W0, let $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{R}_{+}}$ and $\omega\in\Omega$. Then, we have $\omega\in D_{t,f}$ iff $x_{t}(\omega,f)$ has at least two elements. This is equivalent to the fact that $f$ is decreasing on $[0,t)$, $f(t-)=1$, and $P_{u}(\omega)<2$ for all $u\in[0,t]$. Whence the claimed representation of $D_{t,f}$ and the fact that $D_{t,f}\in\mathscr{A}$, because $P$ has continuous paths so that $\max_{u\in[0,t]}P_{u}$ is $\mathscr{A}$-measurable since $P_{u}$ is so for all real (and in particular rational) $u\geq 0$.
Regarding assumption AP.W1, let $w=(\omega,f)\in W$ and $t,u\in\mathbb{R}_{+}$ with $t<u$ such that $x_{t}(w)=x_{u}(w)$. Hence, for all decreasing $g\colon\mathbb{R}_{+}\to\{0,1\}$ with $g|_{[0,t)}=f|_{[0,t)}$ and $(\omega,g)\in W$, we must have
@@
hence $(\omega,g_{a})\in W$. In particular, ()($\ast$)( ∗ ) would imply that $a=g_{a}(t)=f(t)$, for both $a\in\{0,1\}$ which is impossible. Hence, in any possible case we have $x_{t}(w)=\{w\}$, thus assumption AP.W1 is satisfied.
Regarding assumption AP.W2, let $\omega\in\Omega$, $\tilde{f}\in\mathbb{A}^{\mathbb{R}_{+}}$, and $\mathbb{T}^{\prime}\subseteq\mathbb{R}_{+}$ such that $x_{t}(\omega,\tilde{f})\in F$ for all $t\in\mathbb{T}^{\prime}$. Hence, $\tilde{f}$ is decreasing on $[0,\sup\mathbb{T}^{\prime})$ (where, in the context of the order on $\mathbb{R}_{+}$, we have $\sup\emptyset=0$), and if $\tilde{f}|_{[0,\sup\mathbb{T}^{\prime})}$ attains 00, then $P_{u}(\omega)<2$ for all $u\in[0,t_{\tilde{f}}^{\ast}]$. Let $f\colon\mathbb{R}_{+}\to\{0,1\}$ be given by $f|_{[0,\sup\mathbb{T}^{\prime})}=\tilde{f}|_{[0,\sup\mathbb{T}^{\prime})}$ and, for all $u\in[\sup\mathbb{T}^{\prime},\infty)$, $f(u)=f(\sup\mathbb{T}^{\prime}-)$ if $\sup\mathbb{T}^{\prime}>0$ and $f(u)=1$ if $\sup\mathbb{T}^{\prime}=0$. By construction, $f|_{[0,t)}=\tilde{f}|_{[0,t)}$ for all $t\in\mathbb{T}^{\prime}$. Further, $f$ is decreasing. Moreover, if $f$ attains the value 00, then it does so on $[0,\sup\mathbb{T}^{\prime})$ and so does $\tilde{f}$, and $t_{f}^{\ast}=t_{\tilde{f}}^{\ast}$. Hence, $(\omega,f)\in W$. We conclude that assumption AP.W2 is satisfied.
Regarding assumption AP.W3, let $t\in\mathbb{R}_{+}$ and $f,g\in\mathbb{A}^{\mathbb{T}}$ such that $D_{t,f},D_{t,g}\neq\emptyset$. Then, as shown earlier in this proof, $f(t-)=g(t-)=1$ and
@@
Hence, $D_{t,f}\cap D_{t,g}\neq\emptyset$. Thus, assumption AP.W3 is trivially satisfied. ∎
@@ A.3 Section 3
There are exactly two ($\sigma$-)algebras on $\Omega$: the discrete and the trivial one, that is, $\mathcal{P}(\Omega)$ and $\{\Omega,\emptyset\}$. In the present situation, we have $D_{\bm{\mathsf{x}}}=\Omega$ for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$. Hence, by definition, the set of exogenous information structures $\mathscr{F}$ on $(F,\pi,{\bm{\mathsf{X}}})$ is given by all families $\mathscr{F}=(\mathscr{F}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ of $\sigma$-algebras on $\Omega$ such that $\mathscr{F}_{{\bm{\mathsf{x}}}_{0}}\subseteq\mathscr{F}_{{\bm{\mathsf{x}}}_{1}}\cap\mathscr{F}_{{\bm{\mathsf{x}}}_{2}}$. The claim follows easily from this. ∎
@@
There is exactly one ($\sigma$-)algebra on $\{\omega_{2}\}$, namely $\{\{\omega_{2}\},\emptyset\}$. On $\Omega$, there are exactly two $\sigma$-algebras, namely $\mathcal{P}(\Omega)$ and $\{\Omega,\emptyset\}$, as in the preceding proof. Hence, by definition, the set of exogenous information structures $\mathscr{F}^{\prime}$ on $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$ is given by all families $\mathscr{F}^{\prime}=(\mathscr{F}^{\prime}_{{\bm{\mathsf{x}}}^{\prime}})_{{\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}^{\prime}}$ of $\sigma$-algebras on $\Omega$ such that $\mathscr{F}_{{\bm{\mathsf{x}}}^{\prime}_{2}}=\{\{\omega_{2}\},\emptyset\}$ and $\mathscr{F}_{{\bm{\mathsf{x}}}^{\prime}_{0}}\subseteq\mathscr{F}_{{\bm{\mathsf{x}}}^{\prime}_{1}}$. The claim follows easily from this. ∎
@@
Let $\mathscr{F}$, $W$, and $f$ as in the claim. Further, let $t,u\in\mathbb{T}_{f}$ with $t<u$ and $E\in\mathscr{F}_{{\bm{\mathsf{x}}}_{t}(f)}$. Then, ${\bm{\mathsf{x}}}_{t}(f)\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{u}(f)$, whence $E\cap D_{u,f}\in\mathscr{F}_{{\bm{\mathsf{x}}}_{u}(f)}$ by definition 3.24.
The second claim follows from the first because under its hypothesis $\mathscr{F}_{{\bm{\mathsf{x}}}_{t}(f)}$ is a $\sigma$-algebra over $D_{t,f}=\Omega$ for all $t\in\mathbb{T}_{f}$. ∎
@@
Let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$. Then, by construction, $\mathscr{F}_{\bm{\mathsf{x}}}$ is a $\sigma$-algebra over $D_{\bm{\mathsf{x}}}$ contained in $\mathscr{A}$, because $\mathscr{G}_{\mathfrak{t}(x)}$ is a sub-$\sigma$-algebra of $\mathscr{A}$ and all $Y_{{\bm{\mathsf{x}}}^{\prime}}$, ranging over ${\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}$, are $\mathscr{A}$-measurable.

### v2

@@ A.2 Section 2
Take $\mathbb{A}=\{1,2\}$ and $\mathbb{T}=\{0,1\}$. Recall that $\Omega=\{\omega_{1},\omega_{2}\}$, with $\omega_{1}\neq\omega_{2}$, and $\mathscr{E}=\mathcal{P}(\Omega)$.
For the first example, we may set $W=\Omega\times\mathbb{A}^{\mathbb{T}}$. Any $(\omega,f)\in W$ can be identified with the triple $(\omega,f(0),f(1))$, so that $W$ corresponds to the set denoted by $W$ in the example in Subsection 2.3. Under this identification, it is easily verified that $D_{t,f}=\Omega$ for all $f\in\mathbb{A}^{\mathbb{T}}$ and $t\in\mathbb{T}$, ${\bm{\mathsf{x}}}_{0}={\bm{\mathsf{x}}}_{0}(f)$ for all $f\in\mathbb{A}^{\mathbb{T}}$, and ${\bm{\mathsf{x}}}_{k}={\bm{\mathsf{x}}}_{1}(f)$ for all $f\in\mathbb{A}^{\mathbb{T}}$ with $f(0)=k$, for both $k=1,2$. Assumptions AP.SDF0 to AP.SDF3 are readily verified. Under this identification, the action path construction yields exactly the objects $F$, $\pi$, and ${\bm{\mathsf{X}}}$ from the example in Subsection 2.3.
For the second example, let $W^{\prime}$ as in Subsection 2.3. Interpret every triple $(\omega,k,m)\in W^{\prime}$ as the pair $(\omega,f)$ where $f\in\mathbb{A}^{\mathbb{T}}$ is given by $f(0)=k$ and $f(1)=m$. Interpret the pair $(\omega_{1},2)$ as the map $f\colon\mathbb{T}\to\{0,1,2\}$ with $f(0)=2$ and $f(1)=0$. Then, $W\subseteq{\mathbb{A}^{\prime}}^{\mathbb{T}}$ with $\mathbb{A}^{\prime}=\mathbb{A}\cup\{0\}$, where 00 is a placeholder for inaction. Furthermore, $D_{t,f}=\Omega$ for all $(t,f)\in\mathbb{T}\times{\mathbb{A}^{\prime}}^{\mathbb{T}}$ with $t=0$, or $t=1$ and $f(0)=1$; and $D_{1,f}=\{\omega_{2}\}$ for $f\colon\mathbb{T}\to\mathbb{A}^{\prime}$ with $f(0)=2$; for all other pairs $(t,f)\in\mathbb{T}\times{\mathbb{A}^{\prime}}^{\mathbb{T}}$ we have $D_{t,f}=\emptyset$. Moreover, ${\bm{\mathsf{x}}}^{\prime}_{0}={\bm{\mathsf{x}}}_{0}(f)$ for all $f\in{\mathbb{A}^{\prime}}^{\mathbb{T}}$; and ${\bm{\mathsf{x}}}^{\prime}_{1}={\bm{\mathsf{x}}}_{1}(f)$ for all $f\colon\mathbb{T}\to\mathbb{A}^{\prime}$ with $f(0)=1$; and ${\bm{\mathsf{x}}}^{\prime}_{2}={\bm{\mathsf{x}}}_{1}(f)$ for $f\colon\mathbb{T}\to\mathbb{A}^{\prime}$ with $f(0)=2$. Assumptions AP.SDF0 to AP.SDF3 are readily verified. Under these identifications, the action path construction yields exactly the objects $F^{\prime}$, $\pi^{\prime}$, and ${\bm{\mathsf{X}}}^{\prime}$ from the example in Subsection 2.3. ∎
@@
Clearly, for all $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$, $D_{t,f}=\Omega$, whence AP.SDF0.
Regarding AP.SDF1, let $w=(\omega,f)\in W$ and $t,u\in\mathbb{T}$ with $t\neq u$ such that $x_{t}(w)=x_{u}(w)$. Without loss of generality, assume $t<u$. If $\mathbb{A}$ is a singleton, then $\mathbb{A}^{\mathbb{T}}$ is a singleton as well. Hence, $x_{t}(w)=\{w\}$. But $\mathbb{A}$ must be a singleton, since else there would be $g\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,g)\in x_{t}(w)$ and $g(t)\neq f(t)$. Thus, $(\omega,g)\in x_{t}(w)\setminus x_{u}(w)$ – a contradiction.
Regarding AP.SDF2, let $\omega\in\Omega$, $\tilde{f}\in\mathbb{A}^{\mathbb{T}}$, and $\mathbb{T}^{\prime}\subseteq\mathbb{T}$ satisfying $x_{t}(\omega,\tilde{f})\in F$ for all $t\in\mathbb{T}^{\prime}$. Then we already have $(\omega,\tilde{f})\in W$.
AP.SDF3 is clearly satisfied because $D_{t,f}=\Omega$ for all $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$.
(The timing problem):  Clearly, for any $\omega\in\Omega$ there is $f\in\mathbb{A}^{\mathbb{T}}$ with $(\omega,f)\in W$, e.g. the constant map taking only the value $1$ at all times, in all components.
Regarding Assumption AP.SDF0, for all $(t,f)\in\mathbb{A}^{\mathbb{T}}$, $D_{t,f}=\emptyset$ if $f$ is not decreasing on $[0,t)_{\mathbb{T}}$ or if $f(t-)=0$. Else, $D_{t,f}=\Omega$. Hence, AP.SDF0 is satisfied.
Regarding Assumption AP.SDF1, let $w=(\omega,f)\in W$ and $t,u\in\mathbb{T}$ with $t<u$ such that $x_{t}(w)=x_{u}(w)$. Hence, for all decreasing $g\colon\mathbb{T}\to\mathbb{A}$ with $g|_{[0,t)_{\mathbb{T}}}=f|_{[0,t)_{\mathbb{T}}}$ we must have $f(v)=g(v)$ for all $v\in[t,u)_{\mathbb{T}}$. Thus, we must have $f(t-)=0$. Hence, $x_{t}(w)=\{(\omega,f)\}$.
Regarding Assumption AP.SDF2, let $\omega\in\Omega$, $\tilde{f}\in\mathbb{A}^{\mathbb{T}}$, and $\mathbb{T}^{\prime}\subseteq\mathbb{T}$ such that $x_{t}(\omega,\tilde{f})\in F$ for all $t\in\mathbb{T}^{\prime}$. Hence, $\tilde{f}$ is decreasing on $\bigcup_{t\in\mathbb{T}^{\prime}}[0,t)_{\mathbb{T}}$. Hence, there is decreasing $f\colon\mathbb{T}\to\mathbb{A}$ such that $f|_{[0,t)_{\mathbb{T}}}=\tilde{f}|_{[0,t)_{\mathbb{T}}}$ for all $t\in\mathbb{T}^{\prime}$. By construction, $(\omega,f)\in W$.
Regarding Assumption AP.SDF3, it is sufficient to note that $D_{t,f}$ equals $\emptyset$ or $\Omega$, for all $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{T}}$ which has been established earlier.
@@
Regarding Assumption AP.SDF0, let $(t,f)\in\mathbb{T}\times\mathbb{A}^{\mathbb{R}_{+}}$ and $\omega\in\Omega$. Then, we have $\omega\in D_{t,f}$ iff $x_{t}(\omega,f)$ has at least two elements. This is equivalent to the fact that $f$ is decreasing on $[0,t)$, $f(t-)=1$, and $P_{u}(\omega)<2$ for all $u\in[0,t]$. Whence the claimed representation of $D_{t,f}$ and the fact that $D_{t,f}\in\mathscr{E}$, because $P$ has continuous paths so that $\max_{u\in[0,t]}P_{u}$ is $\mathscr{E}$-measurable since $P_{u}$ is so for all real (and in particular rational) $u\geq 0$.
Regarding Assumption AP.SDF1, let $w=(\omega,f)\in W$ and $t,u\in\mathbb{R}_{+}$ with $t<u$ such that $x_{t}(w)=x_{u}(w)$. Hence, for all decreasing $g\colon\mathbb{R}_{+}\to\{0,1\}$ with $g|_{[0,t)}=f|_{[0,t)}$ and $(\omega,g)\in W$, we must have
@@
hence $(\omega,g_{a})\in W$. In particular, ()($\ast$)( ∗ ) would imply that $a=g_{a}(t)=f(t)$, for both $a\in\{0,1\}$ which is impossible. Hence, in any possible case we have $x_{t}(w)=\{w\}$, thus Assumption AP.SDF1 is satisfied.
Regarding Assumption AP.SDF2, let $\omega\in\Omega$, $\tilde{f}\in\mathbb{A}^{\mathbb{R}_{+}}$, and $\mathbb{T}^{\prime}\subseteq\mathbb{R}_{+}$ such that $x_{t}(\omega,\tilde{f})\in F$ for all $t\in\mathbb{T}^{\prime}$. Hence, $\tilde{f}$ is decreasing on $[0,\sup\mathbb{T}^{\prime})$ (where, in the context of the order on $\mathbb{R}_{+}$, we have $\sup\emptyset=0$), and if $\tilde{f}|_{[0,\sup\mathbb{T}^{\prime})}$ attains 00, then $P_{u}(\omega)<2$ for all $u\in[0,t_{\tilde{f}}^{\ast}]$. Let $f\colon\mathbb{R}_{+}\to\{0,1\}$ be given by $f|_{[0,\sup\mathbb{T}^{\prime})}=\tilde{f}|_{[0,\sup\mathbb{T}^{\prime})}$ and, for all $u\in[\sup\mathbb{T}^{\prime},\infty)$, $f(u)=f(\sup\mathbb{T}^{\prime}-)$ if $\sup\mathbb{T}^{\prime}>0$ and $f(u)=1$ if $\sup\mathbb{T}^{\prime}=0$. By construction, $f|_{[0,t)}=\tilde{f}|_{[0,t)}$ for all $t\in\mathbb{T}^{\prime}$. Further, $f$ is decreasing. Moreover, if $f$ attains the value 00, then it does so on $[0,\sup\mathbb{T}^{\prime})$ and so does $\tilde{f}$, and $t_{f}^{\ast}=t_{\tilde{f}}^{\ast}$. Hence, $(\omega,f)\in W$. We conclude that Assumption AP.SDF2 is satisfied.
Regarding Assumption AP.SDF3, let $t\in\mathbb{R}_{+}$ and $f,g\in\mathbb{A}^{\mathbb{T}}$ such that $D_{t,f},D_{t,g}\neq\emptyset$. Then, as shown earlier in this proof, $f(t-)=g(t-)=1$ and
@@
Hence, $D_{t,f}\cap D_{t,g}\neq\emptyset$. Thus, Assumption AP.SDF3 is trivially satisfied. ∎
@@ A.3 Section 3
There are exactly two ($\sigma$-)algebras on $\Omega$: the discrete and the trivial one, that is, $\mathcal{P}(\Omega)$ and $\{\Omega,\emptyset\}$. In the present situation, we have $D_{\bm{\mathsf{x}}}=\Omega$ for all ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$. Hence, by definition, the set of exogenous information structures $\mathscr{F}$ on $(F,\pi,{\bm{\mathsf{X}}})$ admitting recall is given by all families $\mathscr{F}=(\mathscr{F}_{\bm{\mathsf{x}}})_{{\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}}$ of $\sigma$-algebras on $\Omega$ such that $\mathscr{F}_{{\bm{\mathsf{x}}}_{0}}\subseteq\mathscr{F}_{{\bm{\mathsf{x}}}_{1}}\cap\mathscr{F}_{{\bm{\mathsf{x}}}_{2}}$. The claim follows easily from this. ∎
@@
There is exactly one ($\sigma$-)algebra on $\{\omega_{2}\}$, namely $\{\{\omega_{2}\},\emptyset\}$. On $\Omega$, there are exactly two $\sigma$-algebras, namely $\mathcal{P}(\Omega)$ and $\{\Omega,\emptyset\}$, as in the preceding proof. Hence, by definition, the set of exogenous information structures $\mathscr{F}^{\prime}$ on $(F^{\prime},\pi^{\prime},{\bm{\mathsf{X}}}^{\prime})$ admitting recall is given by all families $\mathscr{F}^{\prime}=(\mathscr{F}^{\prime}_{{\bm{\mathsf{x}}}^{\prime}})_{{\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}^{\prime}}$ of $\sigma$-algebras on $\Omega$ such that $\mathscr{F}_{{\bm{\mathsf{x}}}^{\prime}_{2}}=\{\{\omega_{2}\},\emptyset\}$ and $\mathscr{F}_{{\bm{\mathsf{x}}}^{\prime}_{0}}\subseteq\mathscr{F}_{{\bm{\mathsf{x}}}^{\prime}_{1}}$. The claim follows easily from this. ∎
@@
Suppose that the exogenous information structure $\mathscr{F}$ on ${\bm{\mathsf{X}}}$ admits recall and let $f\in\mathbb{A}^{\mathbb{T}}$. Let $t,u\in\mathbb{T}_{f}$ with $t<u$ and $E\in\mathscr{F}_{{\bm{\mathsf{x}}}_{t}(f)}$. Then, ${\bm{\mathsf{x}}}_{t}(f)\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}_{u}(f)$, whence $E\cap D_{u,f}\in\mathscr{F}_{{\bm{\mathsf{x}}}_{u}(f)}$ by Definition 3.27.
The second claim follows from the first because under its hypothesis $\mathscr{F}_{{\bm{\mathsf{x}}}_{t}(f)}$ is a $\sigma$-algebra on $D_{t,f}=\Omega$ for all $t\in\mathbb{T}_{f}$. ∎
@@
Let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$. Then, by construction, $\mathscr{F}_{\bm{\mathsf{x}}}$ is a $\sigma$-algebra on $D_{\bm{\mathsf{x}}}$ contained in $\mathscr{E}$, because $\mathscr{G}_{\mathfrak{t}(x)}$ is a sub-$\sigma$-algebra of $\mathscr{E}$ and all $Y_{{\bm{\mathsf{x}}}^{\prime}}$, ranging over ${\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}$, are $\mathscr{E}$-measurable.